A Bag contains 8 red, 10 white, and 4 blue marbles. Find the probability of obtaining a red or blue marble in a single draw.
I can't seem to understand it because I'm not used to picking two different colors. Not sure what to do.

Answer by ewatrrr(24785)


Hi,
22 Marbles, of which 8 are red, 10 white and 4 blue
P(red 'or' blue in single draw) = 8/22 + 4/22 = 12/22 = 6/11